Stephen has found that the average weight packed gummy candies selected from a normally distributed population is 200g with standard deviation of 8g. Find the point and interval estimates of all gummy candies using 99% confidence level.


1
Expert's answer
2021-05-07T11:43:41-0400

"\bar{x}=200g, \sigma=8g\\

\alpha=0.01,\\

z_{\frac{\alpha}{2}}=z_{0.025}=1.96"


Confidence interval"= \bar{x}\pm z_{0.025}\times \sigma"

"=200\pm 1.96\times 8\\=(200\pm 15.68)\\=(184.32,215.68)"